# SSM Meets Video Diffusion Models: Efficient Video Generation with Structured State Spaces

### v2

@@ 1 Introduction
Recently, state-space models (SSMs) (Gu et al., 2020, 2021, 2022; Smith et al., 2023; Gu & Dao, 2023) have been identified as promising alternatives to attention mechanisms. A pioneering work by Gu et al. (2021) has enabled SSMs to capture long-term dependencies of sequential data, and their model, named the structured state space sequence model (S4), demonstrates superior performance in various benchmarks of sequence modeling. Subsequent studies have also shown that SSMs are effective across various domains, including image and video classification (Nguyen et al., 2022; Knigge et al., 2023; Islam & Bertasius, 2022; Wang et al., 2023), segmentation (Ma et al., 2024), image representation learning (Liu et al., 2024; Zhu et al., 2024), speech generation (Goel et al., 2022), time-series generation (Zhou et al., 2023), language modeling (Mehta et al., 2023; Wang et al., 2022; Dao et al., 2023; Wang et al., 2024), and reinforcement learning (Bar David et al., 2023; Lu et al., 2023; Deng et al., 2023). In contrast to attention mechanisms, SSMs can handle sequential data with linear complexities, so they are expected to overcome the fundamental limitation of attention-based models in many sequence modeling tasks.
However, in the field of video generation, the application of SSMs has not been explored. This is because a methodology to effectively incorporate SSMs into video generation models, such as VDMs, has not been established. In fact, we have empirically observed that a naive approach that directly replace temporal attention layers of VDMs with SSMs works much worse than the original attention-based VDMs. To bridge this gap, in this paper, we investigate an effective approach to incorporate SSMs with video diffusion models. Our proposed model is summarized in Figure 1. A key insight of our investigation is that using bidirectional SSMs is essential to achieve good performance for SSM-based VDMs. In addition, to capture complex nonlinear dynamics of video sequences, adding a multi-layer perceptron (MLP) after the bidirectional SSM is also very effective to improve the generative performance.
@@ 3.1 State Space Models
SSMs are widely used as sequence models that define a mapping from one-dimensional input signals $u(t)$ to one-dimensional output signals $y(t)$. Both continuous and discrete times can be described as follows:
[…]
In this study, we consider discrete-time SSMs. Unlike RNNs but akin to attention mechanisms, linear SSMs are capable of parallel computations through discrete convolutions, facilitated by the Fast Fourier Transform. The memory requirements and parallelizability of the modules managing the time series data are concisely summarized in Table 1.
@@
SSMs on their own are not a complete solution for sequence modeling. They become effective when the discrete-time matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ are derived from a suitable continuous-time state-space model, which enables them to efficiently manage long-range dependencies (Gu et al., 2020). In response, S4 (Gu et al., 2021) aims to learn the continuous-time SSM parameterization of $\mathbf{A},\mathbf{B},\mathbf{C}$ and the discretization rate $\Delta$, utilizing these to derive the requisite discrete-time parameters. Due to the complexity of the original transformation process, Gu et al. (2022) introduced S4D as a simplified diagonalized version of S4 that maintains similar performance in handling long sequences. We chose S4D as a backbone model for our temporal SSM layer.
As delineated in Equations 8 and 9, SSMs are inherently designed to independently handle single-input, single-output (SISO) systems. Therefore, when employing SSMs to manage multidimensional inputs, it is common practice to append a structure to capture dependencies (such as GLU (Dauphin et al., 2017)) between different dimensions of the output after the SSM (Gu et al., 2021, 2022) (see Figure 1 (c)).
@@ 3.2 Temporal SSM Layer for Diffusion Model-based Video Generation
Additionally, we recognize that SSM, while effective, has limitations in integrating information across different dimensions, demonstrated in Equation 8 and Equation 9. To overcome this, we supplement the SSM with a multi-layer perceptron (MLP) after the bidirectional SSM module (Figure 1(b)). Our experiments demonstrate that this addition significantly enhances the model’s performance, proving its importance in the overall architecture.
@@ 5.4 Comparison with Prior SSM Architectures
Table 4: FVD Comparison with SSM layers proposed in prior works across UCF101 16 frames and MineRL Navigate 64 frames. The resolution of both datasets is $32\times 32$ pixels. The dimensionality of the input processed by the temporal layer, $d_{\text{input}}$, is equal to the channel size, $C$, as inferred from the dimensions of the input tensor $\mathbf{X}\in\mathbb{R}^{(B\times H\times W)\times L\times C}$. In the experimental settings described in this table, the channel sizes in the upper layers of the UNet architecture follow a sequential order of 64, 128, 256, and 512, starting from the topmost layer and progressing downwards..

### v3

@@ 1 Introduction
Recently, state-space models (SSMs) (Gu et al., 2020, 2021, 2022; Smith et al., 2023; Gu & Dao, 2023) have been identified as promising alternatives to attention mechanisms. A pioneering work by Gu et al. (2021) has enabled SSMs to capture long-term dependencies of sequential data, and their model, named the structured state space sequence model (S4), demonstrates superior performance in various benchmarks of sequence modeling. In contrast to attention mechanisms, SSMs can handle sequential data with linear complexities, so they are expected to overcome the fundamental limitation of attention-based models in many sequence modeling tasks.
Although subsequent studies have also shown that SSMs are effective across various domains (see 4.2), in the field of video generation, the application of SSMs has not been explored. This is because a methodology to effectively incorporate SSMs into video generation models, such as VDMs, has not been established. In fact, we have empirically observed that a naive approach that directly replace temporal attention layers of VDMs with SSMs works much worse than the original attention-based VDMs. To bridge this gap, in this paper, we investigate an effective approach to incorporate SSMs with video diffusion models. Our proposed model is summarized in Figure 1. A key insight of our investigation is that using bidirectional SSMs is essential to achieve good performance for SSM-based VDMs. In addition, to capture complex nonlinear dynamics of video sequences, adding a multi-layer perceptron (MLP) after the bidirectional SSM is also very effective to improve the generative performance.
@@ 3.1 State Space Models
SSMs are widely used as sequence models that define a mapping from one-dimensional input signals $u_{k}\in\mathbb{R}$ to one-dimensional output signals $y_{k}\in\mathbb{R}$, with $s_{k}\in\mathbb{R}^{D}$ representing the hidden state. Discrete time sequences are processed as follows:
$\mathbf{s}_{k}=\mathbf{\bar{A}}\mathbf{s}_{k-1}+\mathbf{\bar{B}}u_{k},y_{k}=\mathbf{\bar{C}}\mathbf{s}_{k},$ (8)
where $\mathbf{\bar{A}}\in\mathbb{R}^{D\times D}$, $\mathbf{\bar{B}}\in\mathbb{R}^{D\times 1}$, $\mathbf{\bar{C}}\in\mathbb{R}^{1\times D}$. Unlike RNNs but akin to attention mechanisms, linear SSMs are capable of parallel computations through discrete convolutions, facilitated by the Fast Fourier Transform. The memory requirements and parallelizability of the modules managing the time series data are concisely summarized in Table 1.
@@
SSMs on their own are not a complete solution for sequence modeling. They become effective when the discrete-time matrices $\mathbf{\bar{A}},\mathbf{\bar{B}},\mathbf{\bar{C}}$ are derived from a suitable continuous-time state-space model, which enables them to efficiently manage long-range dependencies (Gu et al., 2020). In response, S4 (Gu et al., 2021) aims to learn the continuous-time SSM parameterization of $\mathbf{A},\mathbf{B},\mathbf{C}$ and the discretization rate $\Delta$, utilizing these to derive the requisite discrete-time parameters. Due to the complexity of the original transformation process, Gu et al. (2022) introduced S4D as a simplified diagonalized version of S4 that maintains similar performance in handling long sequences. We chose S4D as a backbone model for our temporal SSM layer.
As delineated in Equations 8, SSMs are inherently designed to independently handle single-input, single-output (SISO) systems. Therefore, when employing SSMs to manage multidimensional inputs, it is common practice to append a structure to capture dependencies (such as GLU (Dauphin et al., 2017)) between different dimensions of the output after the SSM (Gu et al., 2021, 2022) (see Figure 1 (c)).
@@ 3.2 Temporal SSM Layer for Diffusion Model-based Video Generation
Additionally, we recognize that SSM, while effective, has limitations in integrating information across different dimensions, demonstrated in Equation 8. To overcome this, we supplement the SSM with a multi-layer perceptron (MLP) after the bidirectional SSM module (Figure 1(b)). Our experiments demonstrate that this addition significantly enhances the model’s performance, proving its importance in the overall architecture.
@@ 5.2 Benchmark Results
Figure 2: Left: Memory consumption during training with 8 NVIDIA A100 GPUs (40 GB) at a batch size of 8 and resolution of $32\times 32$. Right: Inference time for generating a sample with a single NVIDIA A100 GPU at a resolution of $32\times 32$ and $T=256$.
@@
In Figure 2, we present how training memory consumption and inference time vary with video sequence length for each temporal layer, using $32\times 32$ resolution images. Training memory consumption data is based on a batch size of 8, while inference times reflect sample generation on a single NVIDIA A100 GPU with the number of diffusion time steps $T$ fixed at 256. Notably, for the attention layer, experiments were capped at 275 frames instead of 400 due to memory limitations of the devices, which restrict training of attention-based video diffusion models to a maximum of approximately 275 frames.
@@ 5.4 Comparison with Prior SSM Architectures
Table 4: FVD Comparison with SSM layers proposed in prior works across UCF101 16 frames and MineRL Navigate 64 frames. The resolution of both datasets is $32\times 32$ pixels. The dimensionality of the input processed by the temporal layer, $d_{\text{input}}$, is equal to the channel size, $C$, as inferred from the dimensions of the input tensor $\mathbf{X}\in\mathbb{R}^{(B\times H\times W)\times L\times C}$. In the experimental settings described in this table, the channel sizes in the upper layers of the UNet architecture follow a sequential order of 64, 128, 256, and 512, starting from the topmost layer and progressing downwards.